Courtney wants to mix 30% alcohol with 70% alcohol to form 10 liters of 40% alcohol. How many liters of each should she use?
Answer by stanbon(75887)

Courtney wants to mix 30% alcohol with 70% alcohol to form 10 liters of 40% alcohol. How many liters of each should she use?
-----
Equation:
alc + alc = alc
----
0.30x + 0.70(10-x) = 0.40*10
30x + 70*10 - 70x = 40*10
---
-40x = -30*10
----
x = (3/4)10
---
x = 7.5 liters (amt. of 30% solution needed)
----
10-x = 2.5 liters (amt. of 70% solution needed)
==========================
Cheers,
Stan H.